# Learning to Track Any Points from Human Motion

### v1

@@ 3.2 Filtering Erroneous Tracks with Optical Flow
Method Training DAVIS First DAVIS Strided Kinetics First RoboTAP First [60]
Dataset AJ $<\delta_{avg}^{x}$ OA AJ $<\delta_{avg}^{x}$ OA AJ $<\delta_{avg}^{x}$ OA AJ $<\delta_{avg}^{x}$ OA
Models evaluated at $256\times 256$ resolution.
OmniMotion [63] - - - - $51.7$ $67.5$ $85.3$
Dino-Tracker [59] - - - - $62.3$ $78.2$ $87.5$
TAPNet [14] Kub $33.0$ $48.6$ $78.8$ $38.4$ $53.1$ $82.3$ $38.5$ $54.4$ $80.6$
TAPIR [15] Kub $58.5$ $70.0$ $86.5$ $61.3$ $73.6$ $88.8$ $49.6$ $64.2$ $85.0$ $59.6$ $73.4$ $87.0$
Online TAPIR [60] Kub $56.2$ $70.0$ $86.5$ $51.5$
TAPTR [37] Kub $63.0$ $76.1$ $\underline{91.1}$ $66.3$ $79.2$ $\underline{91.0}$ $49.0$ $64.4$ $85.2$ $60.1$ $75.3$ $86.9$
TAPTRv2 [36] Kub 63.5 $75.9$ 91.4 $66.4$ $78.8$ 91.3 $49.7$ $64.2$ $85.7$
TAPTRv3 [48] Kub $63.2$ $\underline{76.7}$ $91.0$ $\underline{54.5}$ $\underline{67.5}$ 88.2 - - -
BootsTAPIR [16] Kub+15M $61.4$ $74.0$ $88.4$ $66.2$ $78.5$ $90.7$ 54.6 68.4 $\underline{86.5}$ 64.9 80.1 $86.3$
LocoTrack [12] Kub $63.0$ $75.3$ $87.2$ $\underline{67.8}$ $\underline{79.6}$ $89.9$ $52.9$ $66.8$ $85.3$ $62.3$ $76.2$ $\underline{87.1}$
Models evaluated at $384\times 512$ resolution.
PIPs [22] FT [41] $42.2$ $64.8$ $77.7$ $52.4$ $70.0$ $83.6$
CoTracker2 [28] Kub $62.2$ $75.7$ $89.3$ $65.9$ $79.4$ $\underline{89.9}$ $48.8$ $64.5$ $85.8$
Track-On [1] Kub $\underline{65.0}$ $\underline{78.0}$ $\underline{90.8}$ $53.9$ $67.3$ $\underline{87.8}$
CoTracker3 (online) [27] Kub64+15K $64.4$ $76.9$ 91.2 - - - $54.7$ $67.8$ $87.4$
CoTracker3 (offline) [27] Kub64+15K $63.8$ $76.3$ $90.2$ 55.8 $\underline{68.5}$ $\mathbf{88.3}$
LocoTrack [12] Kub $64.8$ $77.4$ $86.2$ $\underline{69.4}$ $\underline{81.3}$ $88.6$ $52.3$ $66.4$ $82.1$
Anthro-LocoTrack (Ours) Kub+1.4K 65.9 78.9 $87.3$ $\mathbf{71.1}$ 82.9 90.3 $\underline{54.8}$ 68.6 $85.3$
@@
3.3 Supervised Training with Pseudo-Labels
Vertex projection, ray casting for visibility prediction, and optical flow-based filtering yield a collection of refined 2D point trajectories $\mathcal{X}_{p,j}$ with associated with visibility labels $v_{p,j}$. These filtered trajectories and their corresponding visibility labels constitute our automatically generated pseudo-label dataset for point tracking. This dataset captures complex human motions and realistic occlusion patterns that are typically absent in synthetic benchmarks. We use this automatically curated dataset to supervise the training of a point tracking model.
4 Experiments
4.1 Experimental Setup
Dataset construction.
We use the pre-trained TokenHMR [17] model for human mesh recovery and SEA-RAFT [65] for obtaining optical flow. To construct our training dataset, we use 1.4K videos from the Let’s Dance dataset [8], which includes a diverse range of dance performances, from solo acts to multi-person scenes. Unlike BootsTAP [16] or CoTracker3 [16], our training videos are entirely non-proprietary. This makes the training and annotation process fully reproducible, which we believe will make a valuable contribution to the point tracking community.
Training details. We fine-tune the pre-trained LocoTrack [12] base model using AnthroTAP-annotated Let’s Dance dataset, which called Anthro-LD. During fine-tuning, we randomly sample data from our dataset and Kubric Panning MOVi-E [15] with equal probability. When training with our dataset, we do not supervise the occlusion prediction for points identified as occluded, as we intentionally filter the dataset for high precision. We augment the filtered dataset using affine transformations, following the approach used in BootsTAP [16]. The hyperparameters are set as follows: $\delta_{\text{cons}}=1.0$, $\tau_{\text{dist}}=1.0$, $\epsilon_{\text{norm}}=2.0$, $\epsilon_{\text{ratio}}=1\times 10^{-6}$, and $\tau_{\text{ratio}}=0.25$.
Optimization is performed using AdamW [40], with the learning rate of $3\times 10^{-4}$ and weight decay set to $1\times 10^{-3}$. The model is trained for 50K steps. We employ a cosine learning rate schedule with a 1000-step warmup, and apply gradient clipping with a maximum norm of 1.0. Training is conducted with a batch size of 1 per GPU, using 4 NVIDIA A6000 GPUs, and converges within 1 day. This is significantly more efficient compared to CoTracker3, which uses 32 A100 80GB GPUs for training, and BootsTAPIR, which uses 256 A100 GPUs. For each batch, we randomly sample 256 tracks and use a resolution of $256\times 256$.
Figure 4: Qualitative Results on the videos from DAVIS Dataset [47]. We present visual comparisons with CoTracker3 [27] and BootsTAPIR [16]. Anthro-LocoTrack (Ours) consistently demonstrates strong performance on highly deformable objects and severe occlusions.
Evaluation protocol. We use the TAP-Vid [14] benchmark to evaluate our model trained on the AnthroTAP-annotated Let’s Dance dataset [8], which we refer to as Anthro-LD. We also evaluate our model on the RoboTAP dataset [60], a real-world robot dataset featuring videos of robotic manipulation tasks. The benchmark uses two evaluation modes based on how query points are selected: the First setup samples the query point from the first visible frame of each trajectory, whereas the Strided setup samples the query point every five timesteps along the trajectory. For evaluation, we use the following metrics, proposed in Doersch et al. [14]: Average Jaccard (AJ), position accuracy ($<\delta_{avg}^{x}$), and Occlusion Accuracy (OA).
4.2 Main Results
Quantitative comparisons. In Table 1, we present quantitative results on the TAP-Vid benchmark using the LocoTrack [12] base model trained with the dataset generated by our pipeline. Our primary comparisons are against BootsTAPIR [16] and CoTracker3 [27] on the DAVIS [46] and Kinetics [31], which consist of real-world videos. BootsTAPIR and CoTracker3 were trained on 15 million and 15,000 real videos respectively, whereas our model is trained on just 1,400 videos.
Despite this significant difference in training data scale, our method achieves notable performance gains. On the DAVIS First subset, our model surpasses BootsTAPIR by 3.3 percentage points on the $<\delta^{x}_{avg}$, demonstrating the data efficiency of our approach. Additionally, it outperforms CoTracker3 by 2.0 percentage points on the same metric.
Qualitative comparisons. In Figure 4, we visually compare the results of CoTracker3 [27], BootsTAPIR [16], and LocoTrack trained with our method. Our model demonstrates more robust tracking performance even in drastic deformation and severe occlusions.
4.3 Ablation and Analysis
Can training on human points generalize to non-human points? In Table 2, we investigate this question by splitting the query points in the TAP-Vid-DAVIS dataset into human and non-human categories, and comparing the performance gain in each group relative to the baseline. We use Mask2Former [10] to identify points on humans. The results show that the performance improvements are significant for both human and non-human points, with an even greater boost observed for non-human points.
Benchmark Method AJ $<\delta^{x}_{avg}$ OA
DAVIS (Human Only) LocoTrack 50.7 42.4 57.8
Anthro-LocoTrack (Ours) 51.2 (+0.5) 43.3 (+0.9) 58.6 (+0.8)
DAVIS (Non-Human) LocoTrack 58.8 73.1 83.9
Anthro-LocoTrack (Ours) 60.8 (+2.0) 75.2 (+2.1) 85.3 (+1.4)
Table 2: Can training on human points generalize to non-human points? To answer this question, we compare the performance by grouping query points on humans and non-human regions separately in TAP-Vid-DAVIS dataset [14]. Our method shows greater improvement on non-human points.
Model Train Strategy Dataset DAVIS
AJ $<\delta^{x}_{avg}$ OA
(I) CoTracker3 Supervised Kubric 63.8 76.3 90.2
(II) CoTracker3 Self-training LD [8] 64.2 (+0.4) 76.5 (+0.2) 89.6 (-0.6)
Table 3: Comparison to self-training method introduced in CoTracker3 using the same video. We compare our pipeline with self-training introduced in Karaev et al. [27]. (I) CoTracker3 model trained on Kubric [21]. (II) fine-tunes the model on the Let’s Dance (LD) dataset using the self-training introduced in the paper. (III) trains CoTracker3 model using our pipeline. Our method shows greater performance boost than self-training introduced in CoTracker3.
Method DAVIS
AJ $<\delta^{x}_{avg}$ OA
Baseline 64.4 76.9 88.6
+ Optical flow based rejection 64.8 77.3 89.1
Table 4: Ablation on optical flow based track rejection.
Training Dataset Data Type Traj Metrics
Complexity Diversity
DriveTrack [3] Real 0.4396 0.0073
PointOdyssey [71] Synthetic 0.5222 0.1597
Kubric [21] Synthetic 0.1772 0.1165
Anthro-LD (Ours) Real 1.2492 0.1008
Table 5: Comparison of track complexity and diversity with existing datasets.
Comparison with self-training strategy of CoTracker3 on identical data. Table 3 compares our approach with the self-training strategy of CoTracker3 [27] on identical data. We fine-tuned the Kubric-pretrained CoTracker3 baseline on the Let’s Dance dataset [8], the same data our pipeline employs, via two methods: the original self-training from CoTracker3, and fine-tuning with pseudo-labels from AnthroTAP. This allowed a direct comparison of both data utilization strategies on the same baseline and data.
Results in Table 3 show self-training by CoTracker3 yielded marginal baseline improvement on Let’s Dance. In contrast, pseudo-labels from AnthroTAP achieved significantly better performance, demonstrating that the quality of training signals from real-world data is important, not merely the use of such data.
Ablation on optical flow based track filtering. In Table 4, we ablate the effect of optical flow-based filtering by comparing it with the baseline that uses trajectories projected from the human mesh and occlusion prediction via ray casting. While the baseline already achieves strong performance, applying trajectory rejection yields a further performance boost.
Comparative analysis on trajectory complexity and diversity. In Table 5, we provide a comparative analysis of prevalent point tracking training datasets, focusing on trajectory complexity and diversity. We quantify trajectory complexity using the mean angular acceleration for each trajectory, thereby accounting for variations in direction. Trajectory diversity is assessed on a per-video basis by computing the mean standard deviation within each video. Anthro-LD, the Let’s Dance dataset [8] annotated using the AnthroTAP pipeline, shows the highest values for complexity and demonstrates greater diversity compared to other real-world datasets, such as DriveTrack [3].
5 Conclusion
In this paper, we presented a novel method for efficiently generating highly complex pseudo-labeled data for point tracking by leveraging the inherent complexities of human motion captured in videos. By fitting SMPL models to real-world human activities, accurately modeling occlusions via ray-casting, and refining trajectories using optical flow consistency, our approach significantly alleviates the bottleneck of manual annotation and synthetic data limitations. The proposed pipeline achieves state-of-the-art results on the TAP-Vid benchmark with orders of magnitude less data than existing methods. Our method thus provides an effective and accessible path forward for robust point tracking in complex, real-world scenarios.
Appendix
Appendix A Additional Ablation and Analysis
Video Length DAVIS
AJ $<\delta^{x}_{avg}$ OA
(I) $24$ 64.7 76.8 89.4
(II) $48$ 64.8 77.3 89.1
(III) $64$ 64.5 77.1 88.9
Table 6: Ablation on training video length.
Frame DAVIS
Dilation AJ $<\delta^{x}_{avg}$ OA
(I) 1 64.6 76.7 89.6
(II) 2 64.8 77.3 89.1
(III) 3 64.2 76.9 89.0
Table 7: Ablation on frame dilation.
Ablation on training video length. In Table 6, we investigate the influence of training video length on model performance. For this analysis, we fine-tuned the LocoTrack [12] base model using our annotated videos, experimenting with distinct clip length. The results demonstrate that utilizing video clips of 48 frames yields the optimal performance in terms of both AJ and $<\delta^{x}_{avg}$.
Ablation on frame dilation. During training, we adjust the frame rate to control the motion speed by sampling video frames with different dilation factors. In Table 7, we test dilations of 1, 2, and 3, corresponding to 1$\times$, 2$\times$, and 3$\times$ faster motion, respectively. When the dilation is too large, the gap between adjacent frames becomes too wide. We found that a dilation factor of 2 yields the best performance in both AJ and $<\delta^{x}_{avg}$.
Figure 5: Qualitative comparison of trajectory complexity. Anthro-LD, a dataset generated from our pipeline, exhibits diverse and complex motions, unlike the rigid and simple trajectories found in DriveTrack [3]
Appendix B Additional Visualization
Qualitative comparison to Drivetrack. In Figure 5, we qualitatively compare the Anthro-LD dataset, generated using videos from the Let’s Dance dataset [8] with our pipeline, to the DriveTrack dataset [3], which is generated using annotated 3D object bounding boxes of cars. Unlike DriveTrack, which features rigid and simple trajectories with similar patterns across all trajectories, our trajectories exhibit highly complex and diverse motion.
Additional Visualization. We visualize videos annotated using our pipeline in Figure 7 and Figure 8. The visualizations highlight the complex motion present in the dataset.
Appendix C Trajectory Complexity and Diversity Analysis
Trajectory Complexity. Trajectory complexity is quantified using the mean angular acceleration magnitude. This metric captures the rate of change in the direction of motion, highlighting non-linear movements and directional variations. For a given visible segment of a trajectory $\mathcal{X}_{p,j}$, consisting of an ordered sequence of points $(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$ (where $\mathbf{x}_{k}$ corresponds to a point $\mathbf{x}_{p,t_{k},j}$ from the trajectory, and $K\geq 3$ is required for at least one angular acceleration value), and assuming a constant time step $\Delta t$ between frames:
$\mathbf{u}_{k}=\mathbf{x}_{k+1}-\mathbf{x}_{k}\quad\text{for }k=0,\ldots,K-1.$
$\theta_{k}=\text{atan2}(\mathbf{u}_{k}^{(1)}\mathbf{u}_{k+1}^{(2)}-\mathbf{u}_{k}^{(2)}\mathbf{u}_{k+1}^{(1)},\mathbf{u}_{k}\cdot\mathbf{u}_{k+1})\quad\text{for }k=0,\ldots,K-2.$
This sequence of angles is then unwrapped to obtain a continuous sequence of angles $\Phi=(\phi_{0},\phi_{1},\ldots,\phi_{K-2})$ by adding multiples of ±$\pm 2\pi$ to eliminate jumps. The sequence of angular velocities $\omega_{k}$ is computed from the unwrapped turning angles:
$\omega_{k}=\frac{\phi_{k}}{\Delta t}\quad\text{for }k=0,\ldots,K-2.$
$\alpha_{k}=\frac{\omega_{k+1}-\omega_{k}}{\Delta t}\quad\text{for }k=0,\ldots,K-3.$
$C_{\text{seg}}=\frac{1}{K-2}\sum_{k=0}^{K-3}|\alpha_{k}|.$
Trajectory Diversity. Trajectory diversity is assessed by computing the mean standard deviation of centered trajectories from a mean centered trajectory. This measures the spatial variability of trajectories in terms of their shape and relative motion, independent of their absolute starting positions. Given a set of $N$ trajectories $\{\mathcal{X}_{1},\mathcal{X}_{2},\ldots,\mathcal{X}_{N}\}$ within a video (where $\mathcal{X}_{i}$ refers to a specific $\mathcal{X}_{p,j}$, and $\mathcal{X}_{i}(t)$ denotes the 2D coordinate $\mathbf{x}_{p,t,j}$ for the $i$-th trajectory at frame $t$):
First, each trajectory $\mathcal{X}_{i}(t)$ is centered by subtracting the coordinates of its first visible point $\mathcal{X}_{i}(t_{i,\text{first\_vis}})$ from all its subsequent visible points. Let this centered trajectory be $\mathcal{X}^{\prime}_{i}(t)$:
$\mathcal{X}^{\prime}_{i}(t)=\mathcal{X}_{i}(t)-\mathcal{X}_{i}(t_{i,\text{first\_vis}})\quad\text{for visible }t\geq t_{i,\text{first\_vis}}.$
$M^{\prime}(t)=\frac{1}{|\mathcal{V}_{t}|}\sum_{i\in\mathcal{V}_{t}}\mathcal{X}^{\prime}_{i}(t).$
$MSD_{i}=\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}||\mathcal{X}^{\prime}_{i}(t)-M^{\prime}(t)||_{2}^{2}.$
@@ Appendix C Trajectory Complexity and Diversity Analysis
$D_{\text{video}}=\text{mean}(SD_{1},SD_{2},\ldots,SD_{N}).$
@@
Appendix D Additional Data Statistics
Statistics on the number of people in training video. In Figure 6, we investigate the number of people per frame in the generated dataset. Most frames contain more than one person (82%), which can lead to complex trajectories due to occlusions between individuals.
Appendix E Limitations and Social Impact

### v2

@@ Appendix C Filtering outlier trajectories with optical flow
Appendix D Trajectory complexity and diversity analysis
Trajectory complexity. Trajectory complexity is quantified using the mean angular acceleration magnitude. This metric captures the rate of change in the direction of motion, highlighting non-linear movements and directional variations. For a given visible segment of a trajectory $\mathcal{X}_{p,j}$, consisting of an ordered sequence of points $(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$ (where $\mathbf{x}_{k}$ corresponds to a point $\mathbf{x}_{p,t_{k},j}$ from the trajectory, requiring at least 4 points with $K\geq 3$ for at least one angular acceleration value), and assuming a constant time step $\Delta t$ between frames:
$\mathbf{u}_{k}=\mathbf{x}_{k+1}-\mathbf{x}_{k}\quad\text{for }k=0,\ldots,K-1.$ (7)
$\theta_{k}=\text{atan2}(\mathbf{u}_{k}^{(1)}\mathbf{u}_{k+1}^{(2)}-\mathbf{u}_{k}^{(2)}\mathbf{u}_{k+1}^{(1)},\mathbf{u}_{k}\cdot\mathbf{u}_{k+1}),$ (8)
for $k=0,\ldots,K-2$. This sequence of angles is then unwrapped to obtain a continuous sequence of angles $\Phi=(\phi_{0},\phi_{1},\ldots,\phi_{K-2})$ by adding multiples of $\pm 2\pi$ to eliminate jumps. The sequence of angular velocities $\omega_{k}$ is computed from the unwrapped turning angles:
$\omega_{k}=\frac{\phi_{k}}{\Delta t}\quad\text{for }k=0,\ldots,K-2.$ (9)
$\alpha_{k}=\frac{\omega_{k+1}-\omega_{k}}{\Delta t}\quad\text{for }k=0,\ldots,K-3.$ (10)
$C_{\text{seg}}=\frac{1}{K-2}\sum_{k=0}^{K-3}|\alpha_{k}|.$ (11)
Trajectory diversity. Trajectory diversity is assessed by computing the mean standard deviation of centered trajectories from a mean centered trajectory. This measures the spatial variability of trajectories in terms of their shape and relative motion, independent of their absolute starting positions.
Consider a set of $N$ trajectories $\{\mathcal{X}_{1},\mathcal{X}_{2},\ldots,\mathcal{X}_{N}\}$ within a single video, where $\mathcal{X}_{i}$ refers to a specific $\mathcal{X}_{p,j}$, and $\mathcal{X}_{i}(t)$ denotes the 2D coordinate $\mathbf{x}_{p,t,j}$ for the $i$-th trajectory at frame $t$. First, each trajectory $\mathcal{X}_{i}(t)$ is centered by subtracting the coordinates of its first visible point $\mathcal{X}_{i}(t_{i,\text{first\_vis}})$ from all its subsequent visible points. Let this centered trajectory be $\mathcal{X}^{\prime}_{i}(t)$:
$\mathcal{X}^{\prime}_{i}(t)=\mathcal{X}_{i}(t)-\mathcal{X}_{i}(t_{i,\text{first\_vis}})\quad\text{for visible }t\geq t_{i,\text{first\_vis}}.$ (12)
$M^{\prime}(t)=\frac{1}{|\mathcal{V}_{t}|}\sum_{i\in\mathcal{V}_{t}}\mathcal{X}^{\prime}_{i}(t).$ (13)
@@ Appendix D Trajectory complexity and diversity analysis
$D_{\text{video}}=\text{mean}(SD_{1},SD_{2},\ldots,SD_{N}).$ (15)
@@
Appendix E Additional data statistics
Statistics on the number of people in training videos. In Figure 6, we investigate the number of people per frame in the generated dataset. Most frames contain more than one person (82%), which can lead to complex trajectories due to occlusions between individuals.
Appendix F Limitations and social impact